OpenStudy (littlebird):

Is this a system of equations? 4x-y=4 -4x+y=-4

ganeshie8 (ganeshie8):

linear system of two dependent equations

OpenStudy (littlebird):

How do I find out if it has a solution?

OpenStudy (igreen):

Graph it or solve it

OpenStudy (igreen):

Multiply -1 to the 2nd equation..what do you get?

OpenStudy (texaschic101):

A system of equations is a collection of 2 or more equations with the same set of unknowns

OpenStudy (littlebird):

In both equations, when solving for y,I got y=4x-4 Does that mean they have no solution, or one unique solution?

OpenStudy (igreen):

\(\sf (-4x + y = -4) \cdot -1 = ~ \! ?\)

OpenStudy (texaschic101):

wrong answer....do what igreen says...it will eventually make sense to you

OpenStudy (littlebird):

4x-y=4

OpenStudy (texaschic101):

oh..I see what you did LittleBird...you see that both equations are equal...they are the same. That means they are the same line......infinite solutions

OpenStudy (littlebird):

I have another similar problem where the lines are parallel. Does that mean they have no solution?

OpenStudy (texaschic101):

Little lesson.. (when finding out how many solutions a system of equations have) In y = mx + b form, the slope is in the m position and the y intercept is in the b position. (when comparing a system of equations).. If the slope and the y intercepts are equal, then it has infinite solutions if the slope is the same, but the y intercepts are different...parallel lines with no solutions. if the slopes and the y intercepts are different, then it has 1 solution

OpenStudy (littlebird):

That is exactly what I needed to know! Thanks!

OpenStudy (texaschic101):

anytime :)